Erin's age is 3 times Warren's. In 4 years she will be twice as old as he will be. How old is each now?
Answer by oscargut(2103)

Let E= actual Erin's age and W=Warren's age so
E=3W and
E+4=2(W+4) so
3W+4=2W+8 so
W=4 and E=12